# Technology Adoption and Network Externalities in Financial Systems: A Spatial-Network Approach

### v1

@@ Abstract
Keywords: Technology adoption, Network externalities, Coordination failures, Lévy processes, Critical mass dynamics, Financial technology
@@ 1 Introduction
This paper develops a unified framework that captures three distinct channels of technology spillovers in financial networks: spatial spillovers reflecting geographic clustering of adopters, network spillovers reflecting adoption by business partners and counterparties, and the interaction between these channels that arises when geographic neighbors are also network-connected. This last channel—absent from existing models—is empirically important in financial technology settings, where institutions form business relationships disproportionately with geographic neighbors. The framework contributes to understanding the externality structure of financial networks in three ways. First, the Adoption Amplification Factor quantifies externalities by measuring how much a shock at one institution affects the entire system beyond the direct effect. Second, the channel decomposition reveals whether externalities flow primarily through network linkages, geographic proximity, or their interaction. Third, the Feynman-Kac representation provides a path-based interpretation: externalities propagate along all possible paths of economic linkage, weighted by probability and discounted by adjustment frictions.
@@
I apply the framework to study SWIFT gpi adoption among Global Systemically Important Banks in 2017. SWIFT gpi represents a major technological upgrade to interbank payment messaging, offering same-day settlement, end-to-end tracking, and confirmation of credit to beneficiary accounts. The adoption pattern provides a natural test of the framework’s predictions: banks with higher amplification factors—those more central in the combined spatial-network structure—should be earlier adopters. The empirical analysis confirms this prediction strongly ($\rho=-0.69$, $p=0.002$), with network-central banks adopting significantly earlier than peripheral banks. Founding members, representing 29 percent of banks in the sample, account for 42 percent of total system amplification, confirming that high-amplification institutions lead adoption. Controlling for network position reveals the role of firm-level characteristics: CEO age delays adoption by 11–15 days per year conditional on firm size and network centrality, consistent with the management literature documenting technology hesitancy among older executives.
The paper proceeds as follows. Section 2 develops the theoretical framework, beginning with the coordinate system and adoption representation, deriving the master equation from three independent economic foundations (heterogeneous agent aggregation, market equilibrium, and cost minimization), presenting the complete master equation with spatial-network interaction, establishing the Feynman-Kac representation and its discrete analog, deriving the Adoption Amplification Factor, demonstrating connections to conventional models through explicit mathematical identification, and extending to Lévy jump-diffusion with state-dependent intensity to capture critical mass dynamics. Section 3 presents Monte Carlo evidence validating the amplification factor as a predictor of technology leadership and demonstrating threshold dynamics. Section 4 applies the framework to SWIFT gpi adoption among Global Systemically Important Banks, presenting the empirical specification, regression results, and graphical evidence. Section 5 discusses extensions including competing technologies and dual externalities. Section 6 concludes.
@@ 1.1 Related Literature
The application to financial technology connects to a growing literature on fintech adoption and banking networks. Buchak et al. (2018) document the growth of fintech lending, while Fuster et al. (2019) study technology’s effect on mortgage lending. The network dimension links to the financial contagion literature: Allen and Gale (2000) and Freixas et al. (2000) develop foundational contagion models, while Acemoglu et al. (2015) characterize how network structure determines whether connections facilitate risk sharing or amplify shocks. The framework developed here analyzes how technology adoption spreads through the same network structures that transmit financial stress, with the key difference that adoption spillovers are typically positive externalities while stress spillovers are negative.
1.2 Related Literature
The application to financial technology connects to a growing literature on fintech adoption and banking networks. Buchak et al. (2018) document the growth of fintech lending, while Fuster et al. (2019) study technology’s effect on mortgage lending. The network dimension links to the financial contagion literature: Allen and Gale (2000) and Freixas et al. (2000) develop foundational contagion models, while Acemoglu et al. (2015) characterize how network structure determines whether connections facilitate risk sharing or amplify shocks. The framework developed here analyzes how technology adoption spreads through the same network structures that transmit financial stress, with the key difference that adoption spillovers are typically positive externalities while stress spillovers are negative.

### v2

@@ Abstract
Keywords: Technology adoption; Network externalities; Coordination failures; Levy processes; Critical mass dynamics; Financial technology
@@ 1 Introduction
This paper develops a unified framework that captures three distinct channels of technology spillovers in financial networks: spatial spillovers reflecting geographic clustering of adopters, network spillovers reflecting adoption by business partners and counterparties, and the interaction between these channels that arises when geographic neighbors are also network-connected. Existing studies of network externalities focus exclusively on network linkages, neglecting spatial spillovers that arise from geographic proximity. However, financial institutions form business relationships disproportionately with geographic neighbors due to information advantages, regulatory similarities, and historical ties. The spatial-network interaction—absent from existing models—captures amplification when both channels operate simultaneously, which is empirically important in financial technology settings. The framework contributes to understanding the externality structure of financial networks in three ways. First, the Adoption Amplification Factor quantifies externalities by measuring how much a shock at one institution affects the entire system beyond the direct effect. Second, the channel decomposition reveals whether externalities flow primarily through network linkages, geographic proximity, or their interaction. Third, the Feynman-Kac representation provides a path-based interpretation: externalities propagate along all possible paths of economic linkage, weighted by probability and discounted by adjustment frictions.
@@
I apply the framework to study SWIFT gpi adoption among Global Systemically Important Banks in 2017. SWIFT gpi represents a major technological upgrade to interbank payment messaging, offering same-day settlement, end-to-end tracking, and confirmation of credit to beneficiary accounts. The adoption pattern provides a natural test of the framework’s predictions: banks with higher amplification factors—those more central in the combined spatial-network structure—should be earlier adopters. The empirical analysis confirms this prediction strongly ($\rho=-0.69$, $p=0.002$), with network-central banks adopting significantly earlier than peripheral banks. Founding members, representing 29 percent of banks in the sample, account for 39 percent of total system amplification, confirming that high-amplification institutions lead adoption. The two-regime dynamics predicted by the Lévy extension are evident in the data: pre-threshold adopters have significantly higher amplification factors than post-threshold adopters (11.81 versus 7.83, $p=0.010$), and the cumulative adoption curve exhibits classic S-curve dynamics. Controlling for network position reveals the role of firm-level characteristics: CEO age delays adoption by 11–15 days per year conditional on firm size and network centrality.
The paper proceeds as follows. Section 2 develops the theoretical framework, beginning with the coordinate system and adoption representation, deriving the master equation from three independent economic foundations (heterogeneous agent aggregation, market equilibrium, and cost minimization), presenting the complete master equation with spatial-network interaction, establishing the Feynman-Kac representation and its discrete analog, deriving the Adoption Amplification Factor, demonstrating connections to conventional models through explicit mathematical identification, and extending to Lévy jump-diffusion with state-dependent intensity to capture critical mass dynamics. Section 3 presents Monte Carlo evidence validating the amplification factor as a predictor of technology leadership and demonstrating threshold dynamics. Section 4 applies the framework to SWIFT gpi adoption among Global Systemically Important Banks, presenting the empirical specification, regression results, and evidence for two-regime dynamics. Section 5 concludes.
@@ 1.1 Related Literature
The application to financial technology connects to a growing literature on fintech adoption and banking networks. Buchak et al. (2018) document the growth of fintech lending, while Fuster et al. (2019) study technology’s effect on mortgage lending. The network dimension links to the financial contagion literature: Allen and Gale (2000) and Freixas et al. (2000) develop foundational contagion models, while Acemoglu et al. (2015) characterize how network structure determines whether connections facilitate risk sharing or amplify shocks. The framework developed here shares the Feynman-Kac foundation and Lévy extension structure with recent work on financial contagion (Kikuchi , 2025), but analyzes positive externalities—technology adoption that benefits counterparties—rather than negative externalities arising from stress transmission. This parallel structure suggests that the same network positions that make institutions systemically important for crisis propagation also make them technology leaders whose adoption decisions cascade broadly through the financial system.